10ml of 0.5M RNH3Cl (kh = 10^-9)+40ml of0.125M KOHFind pH.

Grade:11

1 Answers

Rajat Jain
6 years ago
RNH3Cl+KOH--->RNH3OH+KCl Millimoles:initial:5. 5. 0. 0. Final: 0. 0. 5. 5Kh for RNH3Cl is as good as Ka(acid dissociation constant) for RNH3+.So Kb for RNH3OH will be Kw(Ionic product of water) divided by Ka,since RNH3OH and RNH3+ are conjugate acid-base pair. Now,we have Kb=10^-5So,. RNH3OH --->RNH3+. + OH-Initial Concentration: 5/50. 0. 0Final. 5(1-x)/50 5x./50 5x/50Therefore,. Kb=(5x)(5x)/{5(1-x)*50} :. 10^-5=x^2/10(1-x)Assuming x can be neglected in front of 1 We get x=0.01,which justifies our assumption. Now, [OH-]=5(0.01)/50=10^-3Therefore, pOH=3 and pH=11.Note:Here,contribution of water in concentration of OH- was neglected as [Base]Kb ≥100Kw !!Note-2: In cases,where degree of dissociation of base is negligible and [Base]Kb≥100Kw,[OH-] can be directly written as ([Base]Kb)^0.5!!